VIBRATIONAL ANHARMONICITY AND SCALING THE QUANTUM MECHANICAL MOLECULAR FORCE FIELD

\underline{YURII N. PANCHENKO}, Laboratory of Molecular Spectroscopy, Division of Physical Chemistry, Department of Chemistry, M. V. Lomonosov Moscow State University, Moscow 119899, Russian Federation; VLADIMIR I. PUPYSHEV, Laboratory of Molecular Structure and Quantum Mechanics, Division of Physical Chemistry, Department of Chemistry, M. V. Lomonosov Moscow State University, Moscow 119899, Russian Federation; and CHARLES W. BOCK, Chemistry Department, Philadelphia College of Textiles & Science, Philadelphia, Pennsylvania 19144, U.S.A.

The interrelationship between the scale factors obtained using Pulay’s method ^{a} from the anharmonic and the harmonized vibrational frequencies of a light molecule and its heavy analogue is considered in terms of a Morse potential. The determination of the scale factors from the vibrational frequencies of a light molecule is shown to result in smaller deviations of the calculated and experimental vibrational frequencies of its heavy analogue than those of the reverse procedure. In this context the extent to which Dennison’s rule ^{b} is satisfied is also discussed.

aYu. N. Panchenko, P. Pulay and F. Török, J. Mol. Struct. \underline{34}, 283 (1976); V. I. Pupyshev, Yu. N. Panchenko, Ch. W. Bock and G. Pongor, J. Chem. Phys. \underline{94}, 1247 (1991); Yu. N. Panchenko, G. R. De Maré and V. I. Pupyshev, J. Phys. Chem. \underline{99}, 17544 (1995); Yu. N. Panchenko, Moscow Univ. Chem. Bull. \underline{51} (5), 23 (1996).

bD. M. Dennison, Rev. Mod. Phys. \underline{12}, 175 (1940); G. E. Hansen and D. M. Dennison, J. Chem. Phys. \underline{20}, 313 (1952).